INVESTIGATIONS OF DIMERS AND TRIMERS USING AN UPGRADED INFRARED DIODE LASER SUPERSONIC JET SPECTROMETER

\underline{A. I. MCINTOSH} Z. WANG, R. GARNICA, R. R. LUCCHESE and J. W. BEVAN, Department of Chemistry, Texas A&M University, College Station, Texas 77843-3255.